OpenStudy (anonymous):

Samantha purchased a dining room set for $2,910 using a 12-month deferred payment plan with an interest rate of 22.98%. She did not make any payments during the deferment period. What will Samantha’s monthly payment be if she must pay off the dining room set within two years after the deferment period? $121.25 $160.02 $191.33 $152.24

OpenStudy (anonymous):

b

OpenStudy (anonymous):

How'd you get that? @Courage.

jimthompson5910 (jim_thompson5910):

it's not B how much will the balance be after 12 months?

OpenStudy (anonymous):

It doesn't say?

jimthompson5910 (jim_thompson5910):

use the compound interest formula A = P(1+r/n)^(n*t)

jimthompson5910 (jim_thompson5910):

in this case P = 2910 r = 0.2298 n = 12 t = 1

OpenStudy (anonymous):

I did

jimthompson5910 (jim_thompson5910):

and you got ?

OpenStudy (anonymous):

3653.85

jimthompson5910 (jim_thompson5910):

that's essentially your starting balance when you start to make payments

jimthompson5910 (jim_thompson5910):

from here, you use this formula P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) where, P = monthly payment L = total amount loaned or amortized r = annual interest rate (APR) n = number of times interest is compounded per year (compounding frequency) t = time in years In this case, P = unknown (we're solving for this) L = 3653.85 r = 0.2298 (22.98% = 22.98/100 = 0.2298) n = 12 t = 2

OpenStudy (anonymous):

I got 37.80

OpenStudy (anonymous):

I probably did something wrong but this is all so confusing

OpenStudy (anonymous):

Can you please plug everything in for me?

jimthompson5910 (jim_thompson5910):

that's too low

jimthompson5910 (jim_thompson5910):

P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) P = 3653.85((0.2298/12)*(1 + 0.2298/12)^(12*2))/((1 + 0.2298/12)^(12*2) - 1) P = 191.327925436298 P = 191.33

jimthompson5910 (jim_thompson5910):

I recommend you practice (a lot) with the formula once you get comfortable with it, you can use this calculator to speed things up http://www.bankrate.com/calculators/mortgages/loan-calculator.aspx

OpenStudy (anonymous):

Oh okay I placed everything right my calculator just did something wrong.

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

yw